### v1

# IMUFace: Real-Time, Low-Power, Continuous 3D Facial Reconstruction Through Earphones

Abstract
The potential of facial expression reconstruction technology is significant, with applications in various fields such as human-computer interaction, affective computing, and virtual reality. Recent studies have proposed using ear-worn devices for facial expression reconstruction to address the environmental limitations and privacy concerns associated with traditional camera-based methods. However, these approaches still require improvements in terms of aesthetics and power consumption. This paper introduces a system called IMUFace. It uses inertial measurement units (IMUs) embedded in wireless earphones to detect subtle ear movements caused by facial muscle activities, allowing for covert and low-power facial reconstruction. A user study involving 12 participants was conducted, and a deep learning model named IMUTwinTrans was proposed. The results show that IMUFace can accurately predict users’ facial landmarks with a precision of 2.21 mm, using only five minutes of training data. The predicted landmarks can be utilized to reconstruct a three-dimensional facial model. IMUFace operates at a sampling rate of 30 Hz with a relatively low power consumption of 58 mW. The findings presented in this study demonstrate the real-world applicability of IMUFace and highlight potential directions for further research to facilitate its practical adoption.
Index Terms:
3D facial reconstruction, wearable sensing, mobile computing, low-power
I Introduction
Figure 1: Demonstration of continuous 3D facial reconstruction of IMUFace.
Facial expressions are essential in interpersonal communication because they contain various information, including emotional states [1, 2] and health conditions [3, 4]. Consequently, facial expression reconstruction is widely applied in areas requiring facial analysis. It has become a fundamental technology in fields such as human-computer interaction (HCI), virtual reality (VR), affective computing, and healthcare. For example, facial expression reconstruction can enhance users’ real-time perception of facial expressions and emotions in VR, providing a more immersive experience [5]. In metaverse applications, it can rapidly establish understanding and connections within virtual spaces [6, 7]. Additionally, it can enable silent speech interfaces, facilitating seamless human-computer interaction [8].
Early facial expression reconstruction methods typically relied on manual annotation and geometric models. These traditional methods had low automation and were complex to implement, thus facing significant limitations. With advancements in computer vision and machine learning, automated facial expression reconstruction advances the field. Traditional vision-based methods [9, 10, 11] can accurately locate facial landmarks and generate high-quality facial animations. However, the use of direct camera angles and environmental factors imposes constraints on their efficacy. In order to ensure accuracy, cameras must often face users directly, which limits the applicability of these methods in situations involving varying lighting, resolution, or occlusion conditions. Additionally, the direct capture of users’ faces may give rise to privacy concerns, particularly in contexts where 3D head movements and dynamic environments are involved, as such visual approaches are less suitable in such scenarios.
These limitations have driven a shift toward non-visual facial expression reconstruction methods. Researchers have proposed combining wearable devices and multimodal sensors to enable more flexible and privacy-friendly solutions, such as necklaces [12], hats [13], glasses [14, 15, 16] and ear-mounted devices [17, 18, 19, 20, 21]. Earlier works, like C-Face [22], used miniature ear-mounted cameras to capture facial expressions, offering flexible deployment options. Nonetheless, privacy concerns persist due to camera mobility and proximity to others. Subsequently, BioFace-3D [19] and EarIO [20] introduced non-visual 3D facial reconstruction through ear-mounted devices. However, their designs may cause discomfort by protruding behind the head or slightly touching the face. Later, EARFace [21] embedded acoustic sensors within earplug designs, yielding higher usability scores. However, high power consumption limits acoustic sensors’ viability for prolonged use. Therefore, developing a portable, unobtrusive, low-power system for real-time, continuous facial expression reconstruction remains challenging.
To address these challenges, we present IMUFace, an earphone platform equipped with two IMUs for precise facial reconstruction. These IMUs are highly sensitive, allowing them to capture subtle ear movements caused by facial muscle activity quickly and accurately.
We developed a deep learning model called IMUTwinTrans to achieve 3D facial reconstruction beyond the limitations of cameras. This model, based on the ConvTransformer architecture, is capable of simultaneously leveraging the temporal and frequency features of IMU signals to establish the correspondence between IMU signals and facial landmarks. Additionally, the model adopts a lightweight design to meet the requirements for real-time performance. During testing, the well-trained model can localize facial landmarks directly from IMU signals without visual input. The predicted facial landmarks are then filtered and fitted into a generic 3D head model to generate continuous 3D facial animation. Fig. 1 shows examples of 3D facial reconstruction using the proposed IMUFace system, demonstrating its ability to monitor various parts of the face across different facial motions.
The main contributions of this study are as follows:
  1. 1.
    We presented the first attempt to use IMU sensor signals on wireless earphones for facial landmark tracking to achieve user-friendly, continuous 3D facial expression reconstruction.
  2. 2.
    We designed a lightweight deep learning model, IMUTwinTrans, based on the ConvTransformer architecture, to achieve 3D facial reconstruction by leveraging both temporal and frequency features of IMU signal while meeting real-time performance requirements.
  3. 3.
    We optimized the power consumption of the system, enabling it to operate at a sampling rate of 30 Hz with a power consumption of only 58 mW.
II Related Work
In non-wearable 3D face reconstruction, various algorithms have been developed to address challenges posed by unconstrained photo collections and achieve high-fidelity reconstructions. These methods can be categorized based on their core algorithmic approaches: morphable models, photometric stereo techniques, and deep learning-based solutions.
Morphable models (3DMMs) have been a cornerstone in 3D face reconstruction due to their ability to represent facial geometry and texture in a compact form. Roth et al. [23] introduced an adaptive method for reconstructing 3D face surfaces from unconstrained photo collections using a 3DMM. They employed a coarse-to-fine fitting strategy that allowed the model to adapt even to low-quality image sets, thereby expanding its applicability. This approach improved the quality of reconstructions and handled variations in pose, expression, and illumination more effectively. Roth et al. further advanced this work in [24], where they proposed an algorithm capable of generating accurate 3D face models from diverse 2D images without metadata about cameras or timing. By leveraging a standard field-based Laplace editing technique, they ensured robustness across varied imaging conditions.
Photometric stereo is another key technique used in 3D face reconstruction, especially when dealing with extensive photo collections. Roth et al. [24] utilized this method to capture detailed surface normals from multiple viewpoints, enabling the construction of accurate 3D models. The integration of photometric stereo with landmark constraints provided a powerful means to drive surface reconstruction while maintaining fidelity to the input images. Moreover, Xie et al. [25] explored photometric stereo in the context of smartphone-based reconstruction, demonstrating how calibrated screen lighting can be used to recover fine-grained textures and micro facial features. Their method addresses practical limitations such as hand jitter through the use of face landmarks for alignment and outlier normal elimination techniques.
Deep learning has revolutionized 3D face reconstruction by enabling high-fidelity outputs from single images. Gecer et al. [26] introduced GANFIT, which combines Generative Adversarial Networks (GANs) and Deep Convolutional Neural Networks (DCNNs) to reconstruct facial texture and shape. GANFIT trains a powerful generator of facial texture in UV space and optimizes the parameters under the supervision of pretrained deep identity features. This approach achieves photorealistic and identity-preserving reconstructions with unprecedented detail. Built upon GANFIT, Gecer et al. [27] developed Fast-GANFIT. a faster and more robust version introducing a self-supervised regression-based initialization scheme to expedite the fitting process and enhance stability.
These studies collectively illustrate the evolution of non-wearable 3D face reconstruction algorithms, each offering unique strengths. The versatility of morphable models, the precision of photometric stereo techniques, and the fidelity of deep learning-based solutions all contribute to this advancement. This variety in methodologies highlights the continuous innovation and adaptability present in the field.
II-A 3D Face Reconstruction with Different Models
The 3D Morphable Model (3DMM) [28] serves as the foundational basis for 3D face reconstruction. It represents facial geometry through a linear combination of orthogonal basis functions, with a fixed number of points to depict a face. However, traditional 3DMM approaches require optimizing to solve for relevant coefficients, which is often complex and computationally intensive. To address these challenges, increasing attention has been directed towards leveraging deep learning techniques to directly regress these coefficients, enabling more efficient and precise 3D face reconstruction.
In the realm of deep learning, CNNs have been extensively applied to 3D face reconstruction. For instance, the paper Richardson et al. [29]presents an end-to-end CNN framework that reconstructs detailed facial geometry in a coarse-to-fine manner. Similarly, Tian et al. [30] introduce a CNN-based method for reconstructing 3D faces from two facial images, demonstrating superior performance across various datasets.
Beyond CNNs, Encoder-Decoder architectures have also been employed in 3D face reconstruction. Ji et al. [31] proposes a multi-task encoder-decoder framework based on a Siamese network to learn view-consistent shapes and textures from a single image. Meanwhile, Li et al. [32] enhanced reconstruction accuracy by using a dual-pathway encoder-decoder network to distinguish between different attributes.
Recently, Transformer models have made significant strides in 3D face reconstruction. Chen et al. [33] utilizes a conditional Generative Adversarial Network (cGAN) and a novel mesh transformer for cross-domain face synthesis and reconstruction. It introduces a reprojection consistency loss for self-supervised learning. Another study, Yaermaimaiti et al. [34] integrates ResNet and Transformer modules, leveraging self-attention mechanisms to improve parameter prediction accuracy and significantly enhance the reconstruction results.
II-B Non-wearable 3D Face Reconstruction
The research field of 3D face reconstruction employs various data sources generated by different devices to improve reconstruction quality and applicability. Traditional 3D face reconstruction methods primarily rely on external imaging devices, such as RGB cameras, utilizing extensive photo collections for reconstruction. For instance, Roth et al. [23] present a method capable of reconstructing a 3D face model of an individual from photo collections captured under varied poses and lighting conditions while also extracting albedo information. This approach achieves superior experimental results by fitting a 3D Morphable Model to form a personalized template and developing a novel photometric stereo formulation. Similarly, Roth et al. [24] introduce an algorithm that generates an accurate 3D face model from unconstrained collections of 2D images. This algorithm employs a photometric stereo method along with a novel normal field-based Laplace editing technique, leveraging images captured in diverse poses, and demonstrates both qualitative and quantitative effectiveness.
II-C Wearable 3D Face Reconstruction
In recent years, researchers have begun to explore 3D face reconstruction using wearable miniature cameras, aiming to facilitate face reconstruction in real-life scenarios while minimizing user interference. Chen et al. [22] propose an ear-mounted wearable device that captures subtle changes in facial contours using two miniature cameras and utilizes a deep learning model to continuously output facial feature points, enabling continuous facial expression reconstruction. Additionally, Aoki et al. [35] introduces a glasses-shaped device that captures the user’s facial expressions through two miniature cameras and three mirrors, allowing for the recording and recognition of facial expressions without obstructing daily activities. The system demonstrated high accuracy in emotion recognition, underscoring its applicability to facial image indexing. These studies highlight the expanding potential of wearable devices in 3D face reconstruction, opening new possibilities for practical applications.
The transition from visual methods to multi-modal approaches has led to innovative techniques for monitoring facial expressions and movements using sensors like proximity sensors, EMG, pressure sensors, EOG, and audio signals. Notably, smart glasses and earphones utilizing IMU and acoustic sensors have emerged as effective tools for 3D reconstruction. Wu et al. [19] uses a single-earpiece biosensing system to track facial landmarks and render 3D animations from biosignals without visual input, achieving accuracy comparable to camera-based solutions. Li et al. [16] employ IMUs on smart glasses to estimate action unit intensities for real-time facial reconstruction with low power consumption. Zhang et al. [21] leverages in-ear acoustic sensors to track facial landmarks by detecting changes in the ear canal shape, providing robust and privacy-preserving 3D facial reconstruction. These advancements demonstrate the potential of multi-modal sensors in offering accurate, unobtrusive, and privacy-sensitive facial tracking suitable for diverse applications.
III Background
In this section, we introduce the background information on the following aspects: (1) facial motion and the muscles involved; (2) principle of operation of IMUFace.
III-A Facial Muscles and Expressions
Facial expressions are a crucial element of human communication, providing insight into emotions and intentions that others can perceive and interpret. The human face, equipped with 43 distinct muscles, is capable of producing more than 10,000 different expressions. These muscles, known as mimetic muscles, are organized into three major categories: orbital, nasal, and oral. The orbital muscles control eyelid movements, the nasal group manages movements related to the nose, and the oral muscles facilitate motions of the lips and mouth. Through the interaction of these muscle groups, the human face can depict a wide array of expressions, such as happiness, sadness, surprise, and fear.
The complexity of facial expressions arises from the coordinated activity of these muscle groups, which pull on the facial skin to create the visual cues associated with emotions. For instance, the frontalis muscle raises the eyebrows in surprise, while the temporalis and lateral pterygoid muscles contribute to jaw movements in expressions of amazement or joy.
Figure 2: Overview of the IMUFace system.
III-B Sensing Principle
The relationship between facial expressions and the movement of facial muscles provides a foundation for innovative sensing technologies that capture these expressions indirectly. When facial muscles contract, they alter visible expressions and induce subtle changes in related anatomical structures, such as the ear canal, due to interconnected tissue and bone dynamics.
For example, jaw movements involved in expressions like surprise or joy can affect the mandibular condyle, leading to changes in the shape of the ear canal. Similarly, the temporalis muscle, a major muscle in the head, transmits mechanical artifacts from facial expressions toward the ear area.
An Inertial Measurement Unit (IMU) embedded in an ear-mounted device, like a wireless earbud, can be employed to capture these changes. As facial expressions lead to deformations of the ear canal, the earbud’s position and orientation are altered. The IMU detects these variations through accelerometers and gyroscopes, translating these subtle physical changes into measurable data. Each facial expression has a unique motion pattern, differentiating distinct facial expressions based on the IMU’s output signals. By employing advanced machine learning algorithms, these signals can be analyzed to identify specific facial expressions, offering a practical method to decode them via ear-mounted sensors. This approach opens new possibilities for unobtrusive facial expression recognition, leveraging the natural connection between facial musculature and ear canal dynamics.
Figure 3: Device Prototype: Composed of a pair of earbuds and XLAO nRF52840 Sense MCU.
IV Method
In Fig. 2, we present an overview of our proposed IMUFace system. First, we capture changes in facial expressions using inertial measurement units (IMUs) embedded in earphones, in conjunction with corresponding video data of the human face. The collected data undergoes preprocessing. A quick calibration process is conducted for the IMU data to ensure that the IMUs effectively capture skin movement. Next, we apply the Short-Time Fourier Transform (STFT) to extract frequency domain features from the IMU data, after which we segment this data.
To process the video data, we begin by extracting facial landmarks and normalizing them. These normalized landmarks serve as the ground truth during the training process. Next, we input the preprocessed IMU data into a transformer-based deep learning model. This model uses information from the IMU sensors and analyzes the relationships between consecutive frames to estimate the positions of facial landmarks in the current frame. Finally, we fit these landmarks to the parameters of the FLAME model (Faces Learned with an Articulated Model and Expressions) to create 3D facial reconstruction animations.
IV-A Hardware Prototype Design
We constructed an earphone-based, proof-of-concept prototype capable of capturing IMU signals generated by facial expression movements. It primarily consists of two parts: a tiny sensor board and an earphone frame. We selected the XIAO nRF52840 Sense as the sensor board, with a Bluetooth chip and a 6-axis IMU chip. The Bluetooth chip supports Bluetooth 5.0 for data transmission and directly interfaces with the IMU chip, extracting data using the I2C protocol. The sensor board measures 21.0 mm × 17.5 mm. It provides a compact and efficient solution with low power consumption, meeting our requirements perfectly.
The earphone frame we chose is a standard wireless Bluetooth earphone found in daily life, which can be any standard earphone, ensuring the broad applicability of the system. We stream the 6-axis IMU data from both earphones at a sampling rate of 30 Hz with a 16-bit resolution, ensuring the comprehensiveness of the collected muscle information. The schematic diagram of the device is shown in Fig. 3. First, we attach the battery to the back of the XIAO nRF52840 Sense, then mount it on the earphone. Actually, many modern wireless earphones/earbuds are already equipped with the IMU module inside, which could be easily adapted for deploying the IMUFace approach.
IV-B Data Processing Pipeline
IV-B1 Calibration
Since different users may have different neutral or natural head orientations, we ask users to wear the earbuds and keep their heads naturally upright before starting the video recording. We then record 4 seconds of IMU data from all channels and take the sample mean of these values, referred to as the offset. To compensate for the drift in gyroscope measurements, we adjust the gyroscope readings of each frame by the calculated offset.
IV-B2 Signal Synchronization and Filtering
To ensure synchronization between the two modalities of data streams, we record timestamps for both video data and IMU sensor data. Using the starting timestamp of the video data as a reference, we find the closest corresponding timestamp in the IMU sensor data. Since the sampling rate of the IMU sensor data fluctuates around 30 Hz, we resample the IMU sensor data to 30 Hz. Although our hardware device includes a built-in bandpass filter for reading IMU sensor data, we apply a high-pass filter with a cutoff frequency of 0.1 Hz to improve the signal-to-noise ratio.
IV-B3 Signal Normalization
We utilize the open-source tool OpenFace [36] to extract facial landmarks from facial videos. However, the positions of these landmarks can be affected by factors such as head position, pose variations, and changes in camera angle and placement. Following the approach used in the 300W dataset [37], we standardize the facial landmarks to a consistent reference frame through a series of transformations.
First, we select the nose tip as the origin and shift all landmarks accordingly:
$P_{i}^{\prime}=(x_{i}-x_{\text{nose}},y_{i}-y_{\text{nose}})$ (1)
where $P_{i}=(x_{i},y_{i})$ are the original landmark coordinates, and $(x_{\text{nose}},y_{\text{nose}})$ are the coordinates of the nose tip.
Next, by rotating the coordinates, we align the $x$-axis to the direction from the left outer eye corner to the right outer eye corner. The rotation angle $\theta$ is calculated as:
$\theta=\arctan\left(\frac{y_{\text{right}}-y_{\text{left}}}{x_{\text{right}}-x_{\text{left}}}\right)$ (2)
where $(x_{\text{left}},y_{\text{left}})$ and $(x_{\text{right}},y_{\text{right}})$ represent the coordinates of the left and right outer eye corners, respectively.
The rotated coordinates $P_{i}^{\prime\prime}$ are obtained using the following transformation:
$P_{i}^{\prime\prime}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}x_{i}^{\prime}\\ y_{i}^{\prime}\end{bmatrix}$ (3)
Finally, we normalize the distance between the outer eye corners to 1, ensuring uniform scaling. The normalization factor $d$ is given by:
$d=\sqrt{(x_{\text{right}}^{\prime\prime}-x_{\text{left}}^{\prime\prime})^{2}+(y_{\text{right}}^{\prime\prime}-y_{\text{left}}^{\prime\prime})^{2}}$ (4)
The final normalized coordinates $P_{i}^{\prime\prime\prime}$ are computed as:
$P_{i}^{\prime\prime\prime}=\frac{P_{i}^{\prime\prime}}{d}$ (5)
Fig. 4 shows the facial landmark coordinates before and after alignment.
Figure 4: Illustration of facial landmark alignment.
IV-B4 Short-Time Fourier Transform
Each facial expression movement is associated with unique frequency components. Since the STFT better isolates noise components and enhances the clarity of extracted features from time-series data [38], we use STFT to convert time-domain signals into time-frequency domain spectrograms. The STFT of a signal $x(t)$ is computed as:
[…]
where $w(\tau-t)$ is a window function centered at $t$, $f$ is the frequency, and $X(t,f)$ represents the time-frequency representation of the signal. Frequency-domain features improve the training of our model. We apply STFT to each channel of the IMU data with a window length of 30, a step size of 1, and 32 frequency components. Considering the model complexity, we concatenate six spectrograms (sharing the same time axis) to form a single feature representation input. The concatenation is along the frequency axis, resulting in a final frequency-domain feature dimension of $12\times 17=204$, with a length of IMU signal length.
IV-B5 Data Segmentation
We collect IMU signal data with 12 channels and extract frequency-domain features for each channel. Both the time-domain data and the extracted frequency-domain features are used as inputs to the model. To enable the model to utilize rich temporal relationships when predicting facial landmarks, we input ten continuous frames of data, each with a size of 1 × 216.
Figure 5: Illustration of the model structure.
IV-C Model Architecture and Training Strategy
In this study, we designed a deep-learning model called IMUTwinTrans. This model, based on a ConvTransformer architecture, is capable of simultaneously leveraging the temporal and frequency features of IMU signals to establish the correspondence between IMU signals and facial landmarks, as illustrated in Fig. 5.
IV-C1 CNN-based Feature Extraction
We utilize a CNN to extract robust features from the input data frames. The architecture consists of sequentially stacked 1D convolutional layers with residual connections to enhance feature representation. The network begins with an initial convolutional block followed by three residual blocks that progressively increase the number of channels from 64 to 512 while maintaining a consistent kernel size of 3 and a stride of 1. This configuration allows the network to capture both fine-grained and hierarchical information. To mitigate overfitting, we incorporate dropout layers with a probability of 0.15 after each convolutional layer.
Additionally, each residual block includes shortcut connections to preserve gradient flow during training, improving model stability and convergence. Batch normalization is applied after each convolution to ensure faster training and better generalization. The input frames of size 10×216 are passed through this network, yielding an output of size 10×512, effectively capturing informative features for downstream tasks.
IV-C2 Transformer Structure
We perform value embedding and position encoding on the feature data derived from the CNN output (size 10×512) and feed it into a two-layer Transformer Encoder. The encoder includes a self-attention module within the multi-head attention layers, a multilayer perceptron, and normalization layers. Residual connections are incorporated at both the multi-head attention and multilayer perceptron layers to boost the convergence of the machine learning model. Regarding model parameters, we set $d_{model}$ to 512, $n_{head}$ to 4, $d_{ff}$ to 1024, and the dropout ratio to 0.1 in the encoder.
IV-C3 Loss Function
In the task of facial landmark prediction, standard MSE and MAE loss functions treat all errors equally, regardless of their magnitude. In particular, MSE heavily penalizes large errors while neglecting small ones, resulting in important but less active landmarks (such as pupils compared to lips) not receiving sufficient attention during training. To address this issue, we adopt the Wing Loss function [39], which applies a relatively smooth treatment to small errors while assigning higher weights to large errors. This helps the model predict points closer to the true values more accurately without overly penalizing the model for a few large errors. The loss for each facial landmark is defined as:
Wing $\text{Wing Loss}(x_{i})=\begin{cases}w\log(1+\frac{|x_{i}|}{\epsilon}),&\text{if}|x_{i}|<w\\ |x_{i}|-C,&\text{if}|x_{i}|\geq w\end{cases}$ (7)
where $|x_{i}|$ is the L2 distance between the ground truth and the reconstructed coordinate for the $i$ landmark. $w$ represents the threshold of the small error, which is set to 20 in our case. $\epsilon$ means the curvature in the small error range, and $C=w-wln(1+w/\epsilon)$ which links the linear part and non-linear part together.
IV-C4 Optimization
In addition, the network is trained using the Adam optimizer. The learning rate scheduler uses a Warmup + Cosine Annealing strategy. Initially, during the Warmup phase, the learning rate gradually increases to stabilize gradient updates and prevent divergence. This is followed by the Cosine Annealing phase, where the learning rate smoothly decreases following a cosine curve, allowing for more precise parameter adjustments and reducing overfitting. This strategy accelerates convergence while improving the model’s generalization.
IV-D 3D Reconstruction
To improve the system’s usability and reduce the modeling’s complexity, we seek a compact head model that can easily fit the data while preserving enough details to generate expressive facial animations.
Figure 6: Performance of continuous facial landmark tracking for each participant.
FLAME 3D Head Model. IMUFace uses the FLAME [40] model to represent the face in 3D. FLAME is a statistical head model that learns from approximately 33,000 3D scans. It combines a PCA space for identity shape, simple rotational degrees of freedom for the neck, jaw, and eyeballs, pose-dependent corrective blend shapes, and global expression blend shapes to efficiently represent complex facial geometry and expressions. Mathematically, the model is defined as:
$FLAME(\vec{\beta},\vec{\theta},\vec{\psi}):\mathbb{R}^{|\vec{\beta}|\times|\vec{\theta}|\times|\vec{\psi}|}\to\mathbb{R}^{3N}$ (8)
Optimization-based Shape Reconstruction. We employ a series of optimization steps to obtain a 3D face model that closely resembles the user’s real expressions. First, we perform camera calibration by optimizing the parameters for scale, rotation, and translation to minimize the L2 distance between the 2D projections of the 3D face model’s facial landmarks and the user’s facial landmarks. Second, we optimize the FLAME model parameters (shape, pose, and expression) to further reduce the L2 distance between the 2D projections of the facial landmarks from the 3D face model and the user’s facial landmarks. Ultimately, we can render the user’s 3D face model using the optimal FLAME model parameters.
V Performance Evaluation
To validate the effectiveness of our system, we conducted extensive and detailed user studies to collect training data and evaluate the performance of our system in facial reconstruction. We also performed a series of micro-benchmark tests using the trained model to further ensure the system’s reliability and better understand its capabilities. These tests were designed to examine the system’s performance across various scenarios.
V-A Experimental Methodology
V-A1 Data Collection
In the study, we recruited 12 participants to evaluate IMUFace’s performance. The participants included nine males and three females from the local institution, aged from 18 to 28. The experimental equipment comprises the IMUFace prototype, a Hikvision camera, and a desktop computer. Participants were asked to wear the IMUFace wireless earphones prototype, which was connected to the computer via Bluetooth.
To evaluate the performance of tracking 51 facial landmarks, we focused on seven universal facial expressions of emotion involving happiness, sadness, anger, surprise, fear, disgust, and contempt. Before the experiment, we showed the participants examples of the seven facial expressions and had them repeatedly practice making the corresponding facial expressions while wearing our implemented IMUFace prototype.
Each expression was separated by a neutral facial expression (e.g., a relaxed facial expression). To assist participants with their data collection, seven pictures of the corresponding faces were displayed on a screen for them to imitate. The pace and extent to which each expression was performed were not controlled throughout the experiments. While the users performed the facial motion, IMU data from both earphones were collected at a sampling rate of 30 Hz. To validate the accuracy of IMUFace, the built-in camera recorded videos of the user’s face on the computer at 720p resolution and 30 fps.
Each participant completed 12 sessions of the experiment. In each session, the participants were required to make seven expressions, which were prompted in a random order. Each expression was made twice. Each data collection session lasted for 60 seconds, and the participants took adequate rest between sessions. The earphone prototype was removed and remounted across sessions to validate the robustness to natural changes in sensor positions during daily usage.
V-A2 Experimental Settings
The training was performed on a desktop with an Intel i9-12700K CPU, 32 GB of RAM, and an Nvidia GTX 4090 GPU. We used the Adam optimizer with a learning rate of 0.001. To avoid overfitting issues that may happen in the training process, we added dropouts with a parameter of 0.1 following each RELU activation.
V-A3 Evaluation Metric
The IMUFace system is evaluated primarily by two metrics [19]:
1) Mean Absolute Error (MAE): the absolute difference between the predicted facial landmarks and the ground-truth facial landmarks. It is defined as: $MAE=\frac{1}{N}\times\sum\|g-r\|\times\frac{d_{real}}{d_{norm}}$, where $g$ and $r$ denote the facial landmarks extracted from the ground truth of vision and IMUFace in a normalized coordinate frame. ${d_{real}}$ denotes the actual outer corner distance in millimeters(we measured ${d_{real}}$ for each user who participated in the study). ${d_{norm}}$ denotes the outer corner distance in the normalized coordinate frame. From Section IV-B3, we know that ${d_{norm}}=1$. This provided the MAE in units of millimeters, which is the primary metric used to evaluate IMUFace.
2) Normalized Mean Error (NME): the average error between the predicted facial landmarks and the ground truth facial landmarks, represented by $NME=\frac{1}{N}\times\sum\frac{\|g-r\|}{d_{norm}}\times 100\%$, where $g$ and $r$ denote the facial landmarks extracted from the ground truth of vision and IMUFace in a normalized coordinate frame. ${d_{norm}}$ denotes the outer corner distance in the normalized coordinate frame. From Section IV-B3, we know that ${d_{norm}}=1$.
Figure 7: Domain adaptation.
TABLE I: Comparison result of different Models.
Models Average MAE Average NME
Bi-LSTM [41] 2.57 mm 3.97%
TCN [42] 2.39 mm 3.68%
IF-ConvTransformer [43] 2.27 mm 3.50%
Mamba2 [44] 2.37 mm 3.64%
IMUTwinTrans 2.21 mm 3.40%
V-B User-Dependent Evaluation
For each user’s 12 sessions data, we randomly selected five sessions for training. The rest was used for testing in a randomized cross-validation fashion. Fig. 6(a) illustrates the average MAE & NME and corresponding standard deviations for all the 51 facial landmarks of each participant. Specifically, IMUFace obtains an average of 2.21 mm MAE and 3.40% NME with average standard deviations of 0.99 mm and 0.90%. Among all participants, U4 performed the best, with an MAE of only 1.46, while U8 had the highest MAE at 2.72. Fig. 6(b) depicts the Cumulative Density Function (CDF) of the MAE errors for each individual participant as well as cross-participant cases. 60.0% of the MAEs are less than 2.21, indicating a high level of prediction accuracy. Moreover, different landmarks may exhibit varying error levels due to their differing movement characteristics. Fig. 6(c) visualizes the average MAE for all 51 major landmarks. We found that the reconstructed landmarks on the eyebrows and mouth have relatively more significant errors, while the facial landmarks around the eyes have more minor errors. This is mainly because the eyebrows and mouth change more during facial expression variations.
V-C User-Adaptive Evaluation
Due to the need for improvement in cross-user testing results, we attempted to introduce domain adaptation techniques. First, we trained a model using data from other users, then fine-tuned the model’s linear layers with 2-minute data from a specific user, and finally tested the model using the data from the same user. Fig. 7 compares the performance differences between the user-dependent and domain-adapted models. It can be observed that the performance of the domain-adapted model is significantly improved compared to the user-dependent model, with an average MAE of 2.05 mm for the specific user. This improvement may be attributed to the unique characteristics of each user’s facial muscle movements. Through fine-tuning, the model can learn specific patterns from the target user’s data, thereby enhancing its prediction performance.
V-D Micro-benchmark Tests
V-D1 Model Selection
To verify the effectiveness of our model, we compared it with several deep learning network methods commonly used for IMU data analysis to demonstrate the superiority of our model. We evaluated the performance of different networks under user-dependent experimental settings, and the results are presented in Table I.
Figure 8: Number of earbuds.
Figure 9: Training size vs error.
V-D2 Number of Earphones
To validate the system’s efficacy under single-earphone conditions, we investigated the accuracy when using only a single earphone. As shown in Fig. 8, although using both earphones can provide the best accuracy by integrating sensor data from both ears, the accuracy with individual earphones (left or right) is also very close (2.40 mm, 2.36 mm).
V-D3 Impact of Training Data Size
In practical applications, the amount of data that needs to be collected from users for training can affect user experience. Therefore, it is necessary to explore further the relationship between system performance and the required training data to minimize the burden of data collection on users. Fig. 9 shows the overall system performance when the training data size varies from 1 minute to 5 minutes for each participant. We observe that with 1 minute of training data, the accuracy already reaches 2.66 mm, and with 5 minutes of training data, the accuracy saturates at around 2.21 mm. It is shown that IMUFace can still achieve a reasonable level of accuracy even with a relatively short period of user data collection, which offers a high level of user-friendliness.
V-D4 Latency and Power Consumption
The inference time for the 51 facial landmarks was measured on a single NVIDIA GTX 4090 GPU. Our designed model requires only about 1.26 ms to reconstruct a single frame, which is sufficient for real-time applications.
In addition, we used a power monitoring device to measure the power consumption of IMUFace. Specifically, when the system is idle, that is, the microcontroller unit (MCU) is in sleep mode and does not transmit data via Bluetooth, the average power consumption of IMUFace is 50 mW (13.5 mA @ 3.70 V). When IMUFace actively collects IMU data and transmits it via Bluetooth, the total system power consumption is 58 mW (15.7 mA @ 3.70 V). This indicates that with a 100 mAh lithium polymer battery, IMUFace can continuously record data for up to 6.5 hours, meeting the needs of most applications. We examined the power consumption of recent studies on 3D facial reconstruction and summarized the overall power consumption of the system, including the use of sensing and communication power, in Table II. Our analysis shows that IMUFace consumes the least power among all sensing technologies evaluated, with energy consumption 58.0% lower than that of the next most efficient technology, Bioface-3D [19].
TABLE II: Comparison of System Power Consumption among Related Existing Studies
Studies Sensors Power Consumption
NeckFace [12] Cameras 4 W
C-Face [22] Cameras >4 W
BioFace-3D [19] Biosensors 138 mW
EarIO [20] Acoustics 153.7 mW
EyeEcho [15] Acoustics 167 mW
IMUFace IMUs 58 mW
VI Discussion
Our study aims to use conventional motion sensors, such as accelerometers and gyroscopes, to collect data. These sensors are practical and portable for use in any environment. However, the facial expression movement characteristics of users differ significantly, leading to noticeable variations in model performance across different users. Therefore, the system requires a small amount of data from the target user.
VI-A User-independent Model.
Through the experimental results, we found that the performance of the model after fine-tuning with user data is significantly better than that of the user-independent model. This is because the intensity, response, and sensitivity of facial movement data can vary significantly from person to person, leading to substantial differences between data domains. However, this may reduce usability, as new users must go through a registration phase before using the system. Our future research will explore generating large amounts of virtual IMU data using data from other modalities to train a more generalized user-independent model.
VI-B Effects of Body Movements.
In the system design, we considered the impact of initial head pose changes on the IMU signals. However, repetitive head bounces during activities such as walking can produce significant inertial signal noise [45], which can interfere with the normal operation of the system. Future research can explore the use of targeted filtering methods to minimize the noise impact introduced by such activities.
VII Conclusion
This article proposes and designs IMUFace, a system that reconstructs the 3D facial expressions of users using IMUs that are attached to (and potentially built into) wireless earphones. The system can detect the motion changes associated with facial expressions and achieve user-friendly, low-effort, continuous 3D facial reconstruction. We extracted frequency-domain features from the data and designed a lightweight deep learning model, IMUTwinTrans, based on a ConvTransformer architecture to achieve 3D facial reconstruction by leveraging both temporal and frequency features of the IMU signal. It achieves an average facial landmark tracking accuracy of 2.21 mm across 12 users. The deep learning model inference time is only 1.26 ms, satisfying real-time requirements. Our device consumes only 58 mW of power during operation and can run for 6.25 hours on a 3.7 V, 100 mAh battery. Numerous applications in the areas of affective computing, emotional well-being monitoring, facial expression recognition, AR/VR, accessibility, and user interfaces can be further explored in future work.

### v3

@@
-IMUFace: Real-Time, Low-Power, Continuous 3D Facial Reconstruction Through Earphones
+IMUFace: Towards Always-On 3D Facial Reconstruction via Earphone Inertial Sensing
 Abstract
-The potential of facial expression reconstruction technology is significant, with applications in various fields such as human-computer interaction, affective computing, and virtual reality.
-Recent studies have proposed using ear-worn devices for facial expression reconstruction to address the environmental limitations and privacy concerns associated with traditional camera-based methods.
-However, these approaches still require improvements in terms of aesthetics and power consumption.
+Facial expression reconstruction technology offers considerable potential in areas like human-computer interaction, affective computing, and virtual reality.
+To tackle the privacy challenges and environmental constraints inherent in traditional camera-based systems, researchers have recently introduced ear-worn devices as a viable solution.
+Nevertheless, these methods still demand enhancements, particularly in aspects such as design appeal and energy efficiency, to achieve broader applicability and practical use.
 This paper introduces a system called IMUFace.
@@
 The findings presented in this study demonstrate the real-world applicability of IMUFace and highlight potential directions for further research to facilitate its practical adoption.
-Index Terms:
+keywords:
 3D facial reconstruction, wearable sensing, mobile computing, low-power
-I Introduction
-Figure 1: Demonstration of continuous 3D facial reconstruction of IMUFace.
-Facial expressions are essential in interpersonal communication because they contain various information, including emotional states [1, 2] and health conditions [3, 4].
-Consequently, facial expression reconstruction is widely applied in areas requiring facial analysis.
-It has become a fundamental technology in fields such as human-computer interaction (HCI), virtual reality (VR), affective computing, and healthcare.
-For example, facial expression reconstruction can enhance users’ real-time perception of facial expressions and emotions in VR, providing a more immersive experience [5].
-In metaverse applications, it can rapidly establish understanding and connections within virtual spaces [6, 7].
-Additionally, it can enable silent speech interfaces, facilitating seamless human-computer interaction [8].
-Early facial expression reconstruction methods typically relied on manual annotation and geometric models.
-These traditional methods had low automation and were complex to implement, thus facing significant limitations.
-With advancements in computer vision and machine learning, automated facial expression reconstruction advances the field.
-Traditional vision-based methods [9, 10, 11] can accurately locate facial landmarks and generate high-quality facial animations.
-However, the use of direct camera angles and environmental factors imposes constraints on their efficacy.
-In order to ensure accuracy, cameras must often face users directly, which limits the applicability of these methods in situations involving varying lighting, resolution, or occlusion conditions.
-Additionally, the direct capture of users’ faces may give rise to privacy concerns, particularly in contexts where 3D head movements and dynamic environments are involved, as such visual approaches are less suitable in such scenarios.
-These limitations have driven a shift toward non-visual facial expression reconstruction methods.
-Researchers have proposed combining wearable devices and multimodal sensors to enable more flexible and privacy-friendly solutions, such as necklaces [12], hats [13], glasses [14, 15, 16] and ear-mounted devices [17, 18, 19, 20, 21].
-Earlier works, like C-Face [22], used miniature ear-mounted cameras to capture facial expressions, offering flexible deployment options.
-Nonetheless, privacy concerns persist due to camera mobility and proximity to others.
-Subsequently, BioFace-3D [19] and EarIO [20] introduced non-visual 3D facial reconstruction through ear-mounted devices.
-However, their designs may cause discomfort by protruding behind the head or slightly touching the face.
-Later, EARFace [21] embedded acoustic sensors within earplug designs, yielding higher usability scores.
-However, high power consumption limits acoustic sensors’ viability for prolonged use.
-Therefore, developing a portable, unobtrusive, low-power system for real-time, continuous facial expression reconstruction remains challenging.
+1 Introduction
+Facial expressions play a crucial role in interpersonal communication, as they convey a wide range of information, including emotional states [bassili1979emotion, krumhuber2023role] and physical health conditions [argaud2018facial, roter2006expression].
+As a result, facial expression reconstruction has become an essential tool for applications requiring facial analysis.
+This technology is fundamental to domains such as human-computer interaction (HCI), virtual reality (VR), affective computing, and healthcare.
+For instance, in VR, it improves users’ real-time perception of emotions and facial expressions, creating a more immersive experience [kumari2015facial].
+In metaverse platforms, it enables faster understanding and connection within virtual environments [cha2020real, frith2009role].
+It also supports silent speech interfaces, allowing seamless communication between humans and computers [denby2010silent].
+Traditional methods for reconstructing facial expressions often relied on manual annotation or geometric modeling.
+These early approaches were labor-intensive, difficult to automate, and prone to practical limitations.
+However, with advancements in machine learning and computer vision, automated methods have significantly improved the field.
+Vision-based techniques [cootes2001active, cootes1995active, wang2020deep] enable precise facial landmark detection and high-quality animation generation.
+Nevertheless, their reliance on direct camera views introduces several challenges.
+These methods require users to face the camera under controlled conditions, making them less effective in environments with varying lighting, resolution, or occlusion.
+Additionally, the direct capture of facial imagery raises privacy concerns, particularly in dynamic or sensitive scenarios involving 3D head movements.
+To address these drawbacks, researchers have shifted toward non-visual approaches for facial expression reconstruction.
+Wearable devices combined with multimodal sensors have emerged as a promising solution, offering greater flexibility and enhanced privacy.
+Examples include necklaces [chen2021neckface], hats [bello2023inmyface], glasses [xie2021acoustic, li2024eyeecho, li2024auglasses], and ear-mounted devices [choi2022ppgface, song2022facelistener, wu2021bioface, li2022eario, zhang2023earphone].
+Earlier designs, such as C-Face [chen2020c], utilized miniature cameras mounted on the ear to capture expressions, providing a flexible deployment option.
+However, privacy concerns persisted due to the camera’s proximity and mobility.
+Subsequent innovations like BioFace-3D [wu2021bioface] and EarIO [li2022eario] introduced non-visual 3D facial reconstruction with ear-mounted devices, though their designs sometimes caused discomfort due to protrusions or facial contact.
+More recent developments, such as EARFace [zhang2023earphone], incorporated acoustic sensors into earplug designs, achieving higher usability ratings.
+Nonetheless, the high power consumption of acoustic sensors remains a limitation for extended use.
+Therefore, creating a portable, comfortable, and energy-efficient system capable of real-time facial expression reconstruction continues to be a significant challenge.
 To address these challenges, we present IMUFace, an earphone platform equipped with two IMUs for precise facial reconstruction.
 These IMUs are highly sensitive, allowing them to capture subtle ear movements caused by facial muscle activity quickly and accurately.
+Figure 1: Demonstration of continuous 3D facial reconstruction of IMUFace.
 We developed a deep learning model called IMUTwinTrans to achieve 3D facial reconstruction beyond the limitations of cameras.
@@
 Fig.
-1 shows examples of 3D facial reconstruction using the proposed IMUFace system, demonstrating its ability to monitor various parts of the face across different facial motions.
+1 showcases 3D facial reconstruction by the IMUFace system and, together with the real-time IMU gyroscope signals (X, Y, Z axes), highlights how subtle angular changes from facial muscle movements produce distinct peaks and troughs that correspond to different expressions, confirming both the sensor’s high sensitivity and the physical meaning of the IMU data for deep learning–based facial reconstruction.
 The main contributions of this study are as follows:
@@
     We optimized the power consumption of the system, enabling it to operate at a sampling rate of 30 Hz with a power consumption of only 58 mW.
-II Related Work
-In non-wearable 3D face reconstruction, various algorithms have been developed to address challenges posed by unconstrained photo collections and achieve high-fidelity reconstructions.
-These methods can be categorized based on their core algorithmic approaches: morphable models, photometric stereo techniques, and deep learning-based solutions.
-Morphable models (3DMMs) have been a cornerstone in 3D face reconstruction due to their ability to represent facial geometry and texture in a compact form.
+2 Related Work
+The field of non-wearable 3D face reconstruction has advanced significantly, driven by developments in morphable models, photometric stereo methods, and deep learning approaches.
+These methods address challenges such as variations in pose, illumination, and image quality, enabling high-fidelity reconstruction from unconstrained inputs.
+Morphable Models serve as a foundational tool for compactly representing facial geometry and texture.
 Roth et al.
-[23] introduced an adaptive method for reconstructing 3D face surfaces from unconstrained photo collections using a 3DMM.
-They employed a coarse-to-fine fitting strategy that allowed the model to adapt even to low-quality image sets, thereby expanding its applicability.
-This approach improved the quality of reconstructions and handled variations in pose, expression, and illumination more effectively.
+[roth2016adaptive] introduced an adaptive 3DMM-based method for reconstructing faces from unconstrained photo collections.
+Their approach utilized a coarse-to-fine fitting strategy, enhancing robustness against low-quality images and diverse imaging conditions.
+An earlier work [roth2015unconstrained] proposed an algorithm for generating accurate 3D models from 2D images without relying on metadata, incorporating Laplace editing to improve resilience to pose and illumination variations.
+Photometric Stereo techniques have been pivotal in capturing surface normals from multiple viewpoints, contributing to detailed facial reconstructions.
+For instance, Roth et al.
+[roth2015unconstrained] integrated photometric stereo with landmark constraints to achieve higher reconstruction accuracy.
+Xie et al.
+[xie2021scifi] demonstrated a practical smartphone-based method using calibrated screen lighting, addressing challenges like hand jitter through alignment and outlier removal techniques.
+Deep Learning has transformed 3D face reconstruction, enabling high-fidelity results from single images.
+Gecer et al.
+[gecer2019ganfit] combined Generative Adversarial Networks (GANs) with Deep Convolutional Neural Networks (DCNNs) in GANFIT, achieving photorealistic reconstructions of facial shape and texture.
+Fast-GANFIT [gecer2021fast], their subsequent work, introduced a self-supervised regression-based initialization scheme, improving speed and stability.
+2.1 3D Face Reconstruction with Different Models
+The 3D Morphable Model (3DMM) [blanz2023morphable] remains a cornerstone in 3D face reconstruction.
+While traditional methods optimize coefficients for reconstruction, which is computationally expensive, recent advancements utilize deep learning to directly predict these coefficients, offering greater efficiency and precision.
+Convolutional Neural Networks (CNNs) are widely employed in this field.
+Richardson et al.
+[richardson2017learning] developed an end-to-end CNN framework for reconstructing facial geometry in a coarse-to-fine manner.
+Tian et al.
+[tian2018landmark] presented a CNN-based method for reconstructing 3D faces from two images, demonstrating superior performance on benchmark datasets.
+Encoder-Decoder Architectures have also gained prominence.
+Ji et al.
+[ji2020view] proposed a multi-task encoder-decoder framework based on a Siamese network to ensure view-consistent reconstruction of facial shapes and textures from single images.
+Similarly, Li et al.
+[li2020novel] enhanced accuracy by employing a dual-pathway encoder-decoder network that separates facial attributes.
+Transformer Models have recently achieved remarkable progress.
+Chen et al.
+[chen2022transformer] utilized a conditional Generative Adversarial Network (cGAN) combined with a mesh transformer for cross-domain face synthesis, introducing a reprojection consistency loss for self-supervised learning.
+Yaermaimaiti et al.
+[yaermaimaiti2024research] integrated ResNet and Transformer modules, leveraging self-attention mechanisms to improve parameter prediction and reconstruction quality.
+2.2 Non-wearable 3D Face Reconstruction
+Traditional methods in 3D face reconstruction rely on external imaging devices like RGB cameras to process extensive photo collections.
 Roth et al.
-further advanced this work in [24], where they proposed an algorithm capable of generating accurate 3D face models from diverse 2D images without metadata about cameras or timing.
-By leveraging a standard field-based Laplace editing technique, they ensured robustness across varied imaging conditions.
-Photometric stereo is another key technique used in 3D face reconstruction, especially when dealing with extensive photo collections.
-Roth et al.
-[24] utilized this method to capture detailed surface normals from multiple viewpoints, enabling the construction of accurate 3D models.
-The integration of photometric stereo with landmark constraints provided a powerful means to drive surface reconstruction while maintaining fidelity to the input images.
-Moreover, Xie et al.
-[25] explored photometric stereo in the context of smartphone-based reconstruction, demonstrating how calibrated screen lighting can be used to recover fine-grained textures and micro facial features.
-Their method addresses practical limitations such as hand jitter through the use of face landmarks for alignment and outlier normal elimination techniques.
-Deep learning has revolutionized 3D face reconstruction by enabling high-fidelity outputs from single images.
-Gecer et al.
-[26] introduced GANFIT, which combines Generative Adversarial Networks (GANs) and Deep Convolutional Neural Networks (DCNNs) to reconstruct facial texture and shape.
-GANFIT trains a powerful generator of facial texture in UV space and optimizes the parameters under the supervision of pretrained deep identity features.
-This approach achieves photorealistic and identity-preserving reconstructions with unprecedented detail.
-Built upon GANFIT, Gecer et al.
-[27] developed Fast-GANFIT.
-a faster and more robust version introducing a self-supervised regression-based initialization scheme to expedite the fitting process and enhance stability.
-These studies collectively illustrate the evolution of non-wearable 3D face reconstruction algorithms, each offering unique strengths.
-The versatility of morphable models, the precision of photometric stereo techniques, and the fidelity of deep learning-based solutions all contribute to this advancement.
-This variety in methodologies highlights the continuous innovation and adaptability present in the field.
-II-A 3D Face Reconstruction with Different Models
-The 3D Morphable Model (3DMM) [28] serves as the foundational basis for 3D face reconstruction.
-It represents facial geometry through a linear combination of orthogonal basis functions, with a fixed number of points to depict a face.
-However, traditional 3DMM approaches require optimizing to solve for relevant coefficients, which is often complex and computationally intensive.
-To address these challenges, increasing attention has been directed towards leveraging deep learning techniques to directly regress these coefficients, enabling more efficient and precise 3D face reconstruction.
-In the realm of deep learning, CNNs have been extensively applied to 3D face reconstruction.
-For instance, the paper Richardson et al.
-[29]presents an end-to-end CNN framework that reconstructs detailed facial geometry in a coarse-to-fine manner.
-Similarly, Tian et al.
-[30] introduce a CNN-based method for reconstructing 3D faces from two facial images, demonstrating superior performance across various datasets.
-Beyond CNNs, Encoder-Decoder architectures have also been employed in 3D face reconstruction.
-Ji et al.
-[31] proposes a multi-task encoder-decoder framework based on a Siamese network to learn view-consistent shapes and textures from a single image.
-Meanwhile, Li et al.
-[32] enhanced reconstruction accuracy by using a dual-pathway encoder-decoder network to distinguish between different attributes.
-Recently, Transformer models have made significant strides in 3D face reconstruction.
+[roth2016adaptive] proposed a method for reconstructing personalized 3D face models from photos captured under varying poses and lighting conditions.
+Their approach employed a 3D Morphable Model and a novel photometric stereo formulation to achieve high-quality results.
+Furthermore, Roth et al.
+[roth2015unconstrained] introduced an algorithm that integrates photometric stereo with Laplace editing to handle unconstrained 2D image collections effectively.
+2.3 Wearable 3D Face Reconstruction
+Recent research has focused on wearable devices for 3D face reconstruction, enabling real-time applications with minimal interference.
 Chen et al.
-[33] utilizes a conditional Generative Adversarial Network (cGAN) and a novel mesh transformer for cross-domain face synthesis and reconstruction.
-It introduces a reprojection consistency loss for self-supervised learning.
-Another study, Yaermaimaiti et al.
-[34] integrates ResNet and Transformer modules, leveraging self-attention mechanisms to improve parameter prediction accuracy and significantly enhance the reconstruction results.
-II-B Non-wearable 3D Face Reconstruction
-The research field of 3D face reconstruction employs various data sources generated by different devices to improve reconstruction quality and applicability.
-Traditional 3D face reconstruction methods primarily rely on external imaging devices, such as RGB cameras, utilizing extensive photo collections for reconstruction.
-For instance, Roth et al.
-[23] present a method capable of reconstructing a 3D face model of an individual from photo collections captured under varied poses and lighting conditions while also extracting albedo information.
-This approach achieves superior experimental results by fitting a 3D Morphable Model to form a personalized template and developing a novel photometric stereo formulation.
-Similarly, Roth et al.
-[24] introduce an algorithm that generates an accurate 3D face model from unconstrained collections of 2D images.
-This algorithm employs a photometric stereo method along with a novel normal field-based Laplace editing technique, leveraging images captured in diverse poses, and demonstrates both qualitative and quantitative effectiveness.
-II-C Wearable 3D Face Reconstruction
-In recent years, researchers have begun to explore 3D face reconstruction using wearable miniature cameras, aiming to facilitate face reconstruction in real-life scenarios while minimizing user interference.
-Chen et al.
-[22] propose an ear-mounted wearable device that captures subtle changes in facial contours using two miniature cameras and utilizes a deep learning model to continuously output facial feature points, enabling continuous facial expression reconstruction.
-Additionally, Aoki et al.
-[35] introduces a glasses-shaped device that captures the user’s facial expressions through two miniature cameras and three mirrors, allowing for the recording and recognition of facial expressions without obstructing daily activities.
-The system demonstrated high accuracy in emotion recognition, underscoring its applicability to facial image indexing.
-These studies highlight the expanding potential of wearable devices in 3D face reconstruction, opening new possibilities for practical applications.
-The transition from visual methods to multi-modal approaches has led to innovative techniques for monitoring facial expressions and movements using sensors like proximity sensors, EMG, pressure sensors, EOG, and audio signals.
-Notably, smart glasses and earphones utilizing IMU and acoustic sensors have emerged as effective tools for 3D reconstruction.
+[chen2020c] designed an ear-mounted device equipped with two miniature cameras to capture subtle facial contour changes.
+Their deep learning model continuously outputs facial feature points, enabling real-time expression reconstruction.
+Aoki et al.
+[aoki2021facerecglasses] developed a glasses-shaped device featuring two cameras and three mirrors to capture facial expressions unobtrusively, achieving high accuracy in emotion recognition.
+Multimodal Approaches have further enhanced wearable devices.
 Wu et al.
-[19] uses a single-earpiece biosensing system to track facial landmarks and render 3D animations from biosignals without visual input, achieving accuracy comparable to camera-based solutions.
+[wu2021bioface] introduced a single-earpiece biosensing system that tracks facial landmarks and generates 3D animations from biosignals, achieving performance comparable to camera-based methods.
 Li et al.
-[16] employ IMUs on smart glasses to estimate action unit intensities for real-time facial reconstruction with low power consumption.
+[li2024auglasses] utilized IMUs in smart glasses to estimate action unit intensities, enabling real-time, low-power facial reconstruction.
 Zhang et al.
-[21] leverages in-ear acoustic sensors to track facial landmarks by detecting changes in the ear canal shape, providing robust and privacy-preserving 3D facial reconstruction.
-These advancements demonstrate the potential of multi-modal sensors in offering accurate, unobtrusive, and privacy-sensitive facial tracking suitable for diverse applications.
-III Background
-In this section, we introduce the background information on the following aspects: (1) facial motion and the muscles involved; (2) principle of operation of IMUFace.
-III-A Facial Muscles and Expressions
-Facial expressions are a crucial element of human communication, providing insight into emotions and intentions that others can perceive and interpret.
-The human face, equipped with 43 distinct muscles, is capable of producing more than 10,000 different expressions.
-These muscles, known as mimetic muscles, are organized into three major categories: orbital, nasal, and oral.
-The orbital muscles control eyelid movements, the nasal group manages movements related to the nose, and the oral muscles facilitate motions of the lips and mouth.
-Through the interaction of these muscle groups, the human face can depict a wide array of expressions, such as happiness, sadness, surprise, and fear.
-The complexity of facial expressions arises from the coordinated activity of these muscle groups, which pull on the facial skin to create the visual cues associated with emotions.
-For instance, the frontalis muscle raises the eyebrows in surprise, while the temporalis and lateral pterygoid muscles contribute to jaw movements in expressions of amazement or joy.
+[zhang2023earphone] employed in-ear acoustic sensors to track facial landmarks by detecting ear canal shape changes, providing a privacy-preserving solution for 3D facial reconstruction.
+3 Background
+This section provides an overview of the foundational concepts, covering: (1) the role of facial muscles in generating expressions and (2) the operational principle of IMUFace.
 Figure 2: Overview of the IMUFace system.
-III-B Sensing Principle
-The relationship between facial expressions and the movement of facial muscles provides a foundation for innovative sensing technologies that capture these expressions indirectly.
-When facial muscles contract, they alter visible expressions and induce subtle changes in related anatomical structures, such as the ear canal, due to interconnected tissue and bone dynamics.
-For example, jaw movements involved in expressions like surprise or joy can affect the mandibular condyle, leading to changes in the shape of the ear canal.
-Similarly, the temporalis muscle, a major muscle in the head, transmits mechanical artifacts from facial expressions toward the ear area.
-An Inertial Measurement Unit (IMU) embedded in an ear-mounted device, like a wireless earbud, can be employed to capture these changes.
-As facial expressions lead to deformations of the ear canal, the earbud’s position and orientation are altered.
-The IMU detects these variations through accelerometers and gyroscopes, translating these subtle physical changes into measurable data.
-Each facial expression has a unique motion pattern, differentiating distinct facial expressions based on the IMU’s output signals.
-By employing advanced machine learning algorithms, these signals can be analyzed to identify specific facial expressions, offering a practical method to decode them via ear-mounted sensors.
-This approach opens new possibilities for unobtrusive facial expression recognition, leveraging the natural connection between facial musculature and ear canal dynamics.
-Figure 3: Device Prototype: Composed of a pair of earbuds and XLAO nRF52840 Sense MCU.
-IV Method
-In Fig.
-2, we present an overview of our proposed IMUFace system.
-First, we capture changes in facial expressions using inertial measurement units (IMUs) embedded in earphones, in conjunction with corresponding video data of the human face.
-The collected data undergoes preprocessing.
-A quick calibration process is conducted for the IMU data to ensure that the IMUs effectively capture skin movement.
-Next, we apply the Short-Time Fourier Transform (STFT) to extract frequency domain features from the IMU data, after which we segment this data.
-To process the video data, we begin by extracting facial landmarks and normalizing them.
-These normalized landmarks serve as the ground truth during the training process.
-Next, we input the preprocessed IMU data into a transformer-based deep learning model.
-This model uses information from the IMU sensors and analyzes the relationships between consecutive frames to estimate the positions of facial landmarks in the current frame.
-Finally, we fit these landmarks to the parameters of the FLAME model (Faces Learned with an Articulated Model and Expressions) to create 3D facial reconstruction animations.
-IV-A Hardware Prototype Design
+3.1 Facial Muscles and Expressions
+Facial expressions play an essential role in human communication by conveying emotions and intentions that can be observed and interpreted by others.
+The human face contains 43 distinct mimetic muscles, which collectively enable the production of over 10,000 unique expressions.
+These muscles are categorized into three primary groups: orbital, nasal, and oral.
+The orbital muscles are responsible for eyelid movements, the nasal muscles manage actions around the nose, and the oral muscles control the lips and mouth.
+The coordination of these groups enables the depiction of a broad range of emotions, such as happiness, sadness, surprise, and fear.
+The intricate nature of facial expressions stems from the synchronized activity of these muscle groups, which exert forces on the skin to form visible emotional cues.
+For instance, the frontalis muscle elevates the eyebrows during expressions of surprise, while jaw movements in amazement or joy involve the temporalis and lateral pterygoid muscles.
+3.2 Sensing Principle
+The interplay between facial muscle activity and associated anatomical changes provides the basis for innovative sensing methods that indirectly capture facial expressions.
+Muscle contractions during expressions not only generate visible changes but also induce subtle structural alterations in nearby areas, such as the ear canal, due to the dynamics of interconnected tissues and bones.
+For example, jaw movements that occur during expressions like joy or surprise influence the mandibular condyle, causing changes in the ear canal’s shape.
+Similarly, the temporalis muscle, a key component of the head’s musculature, transmits mechanical effects of facial expressions to the ear region.
+By embedding an Inertial Measurement Unit (IMU) in an ear-mounted device, such as a wireless earbud, these changes can be detected.
+As facial expressions deform the ear canal, they alter the earbud’s position and orientation.
+The IMU captures these variations using accelerometers and gyroscopes, converting them into measurable signals.
+Each facial expression produces a distinct motion pattern, allowing the IMU to differentiate expressions based on the recorded data.
+Advanced machine learning algorithms process these signals to identify specific expressions, enabling unobtrusive and practical facial expression recognition.
+This approach capitalizes on the natural linkage between facial musculature and ear canal dynamics, offering a novel solution for wearable expression recognition systems.
+4 Method
+In Fig. 2, we present an overview of the architecture of our proposed IMUFace system, which is divided into two distinct phases: a training phase and a testing phase.
+The training phase aims to teach a deep learning model to map IMU signals to facial movements using two synchronized data streams: video and IMU.
+The video stream provides ground truth by detecting and aligning faces in each frame, from which 2D facial landmarks are extracted.
+Meanwhile, the IMU stream, collected from earphone-embedded sensors, captures subtle facial muscle movements.
+This data undergoes a three-step preprocessing pipeline: a quick calibration to adapt to individual facial structures, a Short-Time Fourier Transform (STFT) to convert signals into the frequency domain, and windowed segmentation for temporal analysis.
+The processed IMU data is then fed into our transformer-based model, IMUTwinTrans, which predicts facial landmarks.
+A loss function measures the difference between the predicted and ground-truth landmarks, enabling the model to optimize its parameters through backpropagation.
+Once training is complete, the system enters the testing phase for inference.
+In this phase, only the IMU data stream is required.
+The incoming IMU data undergoes the same preprocessing steps as in the training phase (calibration, STFT, and segmentation).
+This processed data is then fed into the fully trained model.
+The model, having learned the correlation between IMU signals and facial expressions, directly outputs the estimated facial landmarks in real-time.
+As the final step, these predicted landmarks are fitted to the parameters of the FLAME model[li2017learning] to reconstruct and animate a realistic 3D facial avatar.
+Figure 3: Device Prototype: Composed of a pair of earbuds and XIAO nRF52840 Sense MCU.
+4.1 Hardware Prototype Design
 We constructed an earphone-based, proof-of-concept prototype capable of capturing IMU signals generated by facial expression movements.
@@
 Actually, many modern wireless earphones/earbuds are already equipped with the IMU module inside, which could be easily adapted for deploying the IMUFace approach.
-IV-B Data Processing Pipeline
-IV-B1 Calibration
+4.2 Data Processing Pipeline
+4.2.1 Calibration
 Since different users may have different neutral or natural head orientations, we ask users to wear the earbuds and keep their heads naturally upright before starting the video recording.
@@
 To compensate for the drift in gyroscope measurements, we adjust the gyroscope readings of each frame by the calculated offset.
-IV-B2 Signal Synchronization and Filtering
+4.2.2 Signal Synchronization and Filtering
 To ensure synchronization between the two modalities of data streams, we record timestamps for both video data and IMU sensor data.
@@
 Although our hardware device includes a built-in bandpass filter for reading IMU sensor data, we apply a high-pass filter with a cutoff frequency of 0.1 Hz to improve the signal-to-noise ratio.
-IV-B3 Signal Normalization
+Figure 4: Illustration of facial landmark alignment. (a) Before landmark alignment. (b) After landmark alignment.
+4.2.3 Signal Normalization
 We utilize the open-source tool OpenFace [8373812] to extract facial landmarks from facial videos.
@@
 Following the approach used in the 300W dataset [6755925], we standardize the facial landmarks to a consistent reference frame through a series of transformations.
-First, we select the nose tip as the origin and shift all landmarks accordingly:
+First, we select the nose tip as the origin and shift all landmarks accordingly, as defined in Eq. (1):
 $P_{i}^{\prime}=(x_{i}-x_{\text{nose}},y_{i}-y_{\text{nose}})$ (1)
-where $P_{i}=(x_{i},y_{i})$ are the original landmark coordinates, and $(x_{\text{nose}},y_{\text{nose}})$ are the coordinates of the nose tip.
-Next, by rotating the coordinates, we align the $x$-axis to the direction from the left outer eye corner to the right outer eye corner. The rotation angle $\theta$ is calculated as:
+where $P_{i}=(x_{i},y_{i})$ denotes the original landmark coordinates, and $(x_{\text{nose}},y_{\text{nose}})$ represents the position of the nose tip.
+Next, we perform a coordinate rotation to align the $x$-axis with the direction from the left outer eye corner to the right outer eye corner. The rotation angle $\theta$ is computed using Eq. (2):
 $\theta=\arctan\left(\frac{y_{\text{right}}-y_{\text{left}}}{x_{\text{right}}-x_{\text{left}}}\right)$ (2)
-where $(x_{\text{left}},y_{\text{left}})$ and $(x_{\text{right}},y_{\text{right}})$ represent the coordinates of the left and right outer eye corners, respectively.
-The rotated coordinates $P_{i}^{\prime\prime}$ are obtained using the following transformation:
+where $(x_{\text{left}},y_{\text{left}})$ and $(x_{\text{right}},y_{\text{right}})$ correspond to the coordinates of the outer corners of the left and right eyes, respectively.
+The rotated landmark coordinates $P_{i}^{\prime\prime}$ are then obtained by applying the transformation in Eq. (3):
 $P_{i}^{\prime\prime}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}x_{i}^{\prime}\\ y_{i}^{\prime}\end{bmatrix}$ (3)
-Finally, we normalize the distance between the outer eye corners to 1, ensuring uniform scaling. The normalization factor $d$ is given by:
+To ensure uniform scaling, we normalize the distance between the two outer eye corners to 1. The normalization factor $d$ is calculated according to Eq. (4):
 $d=\sqrt{(x_{\text{right}}^{\prime\prime}-x_{\text{left}}^{\prime\prime})^{2}+(y_{\text{right}}^{\prime\prime}-y_{\text{left}}^{\prime\prime})^{2}}$ (4)
-The final normalized coordinates $P_{i}^{\prime\prime\prime}$ are computed as:
+Finally, the normalized landmark coordinates $P_{i}^{\prime\prime\prime}$ are obtained by applying Eq. (5):
 $P_{i}^{\prime\prime\prime}=\frac{P_{i}^{\prime\prime}}{d}$ (5)
 Fig. 4 shows the facial landmark coordinates before and after alignment.
-Figure 4: Illustration of facial landmark alignment.
-IV-B4 Short-Time Fourier Transform
+4.2.4 Short-Time Fourier Transform
 Each facial expression movement is associated with unique frequency components.
-Since the STFT better isolates noise components and enhances the clarity of extracted features from time-series data [38], we use STFT to convert time-domain signals into time-frequency domain spectrograms.
-The STFT of a signal $x(t)$ is computed as:
-[…]
+Since the STFT better isolates noise components and enhances the clarity of extracted features from time-series data [zhang2022force], we use STFT to convert time-domain signals into time-frequency domain spectrograms.
+The STFT of a signal $x(t)$ is computed as Eq.
+(6):
+$X(t,f)=\int_{-\infty}^{\infty}x(\tau)w(\tau-t)e^{-j2\pi f\tau}\,d\tau$ (6)
 where $w(\tau-t)$ is a window function centered at $t$, $f$ is the frequency, and $X(t,f)$ represents the time-frequency representation of the signal.
@@
 Considering the model complexity, we concatenate six spectrograms (sharing the same time axis) to form a single feature representation input.
-The concatenation is along the frequency axis, resulting in a final frequency-domain feature dimension of $12\times 17=204$, with a length of IMU signal length.
-IV-B5 Data Segmentation
+The concatenation is along the frequency axis, resulting in a final frequency-domain feature dimension of $12\times 17=204$, with a length of IMU signal.
+Figure 5: Illustration of the model structure.
+4.2.5 Data Segmentation
 We collect IMU signal data with 12 channels and extract frequency-domain features for each channel.
 Both the time-domain data and the extracted frequency-domain features are used as inputs to the model.
-To enable the model to utilize rich temporal relationships when predicting facial landmarks, we input ten continuous frames of data, each with a size of 1 × 216.
-Figure 5: Illustration of the model structure.
-IV-C Model Architecture and Training Strategy
+To help the model capture temporal relationships when predicting facial landmarks, we use ten consecutive frames of processed IMU data as one input segment.
+4.3 Model Architecture and Training Strategy
 In this study, we designed a deep-learning model called IMUTwinTrans.
-This model, based on a ConvTransformer architecture, is capable of simultaneously leveraging the temporal and frequency features of IMU signals to establish the correspondence between IMU signals and facial landmarks, as illustrated in Fig.
+This model, based on a ConvTransformer architecture, leverages both temporal and frequency features of IMU signals to establish the correspondence between IMU data and facial landmarks, as illustrated in Fig.
 5.
-IV-C1 CNN-based Feature Extraction
+The model input consists of 10 consecutive frames of IMU data that include both time-domain and frequency-domain features.
+Specifically, the time-domain signal for each frame is arranged as a $30\times 12$ matrix.
+The frequency-domain features are obtained via a Short-Time Fourier Transform (STFT), resulting in a $30\times 204$ matrix.
+By concatenating the time-domain and frequency-domain features, each frame forms a combined feature matrix of $30\times 216$.
+The input data are first processed by a Convolutional Neural Network (CNN) for feature extraction, followed by a Transformer encoder to capture temporal dependencies.
+Finally, a linear projection layer outputs the predicted coordinates of 51 facial landmarks, represented as a $1\times 102$ vector.
+4.3.1 CNN-based Feature Extraction
 We utilize a CNN to extract robust features from the input data frames.
@@
 The input frames of size 10×216 are passed through this network, yielding an output of size 10×512, effectively capturing informative features for downstream tasks.
-IV-C2 Transformer Structure
-We perform value embedding and position encoding on the feature data derived from the CNN output (size 10×512) and feed it into a two-layer Transformer Encoder.
+4.3.2 Transformer Structure
+We perform value embedding and position encoding on the feature data derived from the CNN output (size $10\times 512$) and feed it into a two-layer Transformer Encoder.
 The encoder includes a self-attention module within the multi-head attention layers, a multilayer perceptron, and normalization layers.
@@
 Regarding model parameters, we set $d_{model}$ to 512, $n_{head}$ to 4, $d_{ff}$ to 1024, and the dropout ratio to 0.1 in the encoder.
-IV-C3 Loss Function
+4.3.3 Loss Function
 In the task of facial landmark prediction, standard MSE and MAE loss functions treat all errors equally, regardless of their magnitude.
 In particular, MSE heavily penalizes large errors while neglecting small ones, resulting in important but less active landmarks (such as pupils compared to lips) not receiving sufficient attention during training.
-To address this issue, we adopt the Wing Loss function [39], which applies a relatively smooth treatment to small errors while assigning higher weights to large errors.
+To address this issue, we adopt the Wing Loss function [feng2018wing], which applies a relatively smooth treatment to small errors while assigning higher weights to large errors.
 This helps the model predict points closer to the true values more accurately without overly penalizing the model for a few large errors.
-The loss for each facial landmark is defined as:
-Wing $\text{Wing Loss}(x_{i})=\begin{cases}w\log(1+\frac{|x_{i}|}{\epsilon}),&\text{if}|x_{i}|<w\\ |x_{i}|-C,&\text{if}|x_{i}|\geq w\end{cases}$ (7)
+The loss for each facial landmark is defined as (Eq.
+(7)):
+$\text{Wing Loss}(x_{i})=\begin{cases}w\log(1+\frac{|x_{i}|}{\epsilon}),&\text{if}|x_{i}|<w\\ |x_{i}|-C,&\text{if}|x_{i}|\geq w\end{cases}$ (7)
 where $|x_{i}|$ is the L2 distance between the ground truth and the reconstructed coordinate for the $i$ landmark.
@@
 $\epsilon$ means the curvature in the small error range, and $C=w-wln(1+w/\epsilon)$ which links the linear part and non-linear part together.
-IV-C4 Optimization
+4.3.4 Optimization
 In addition, the network is trained using the Adam optimizer.
@@
 This strategy accelerates convergence while improving the model’s generalization.
-IV-D 3D Reconstruction
-To improve the system’s usability and reduce the modeling’s complexity, we seek a compact head model that can easily fit the data while preserving enough details to generate expressive facial animations.
-Figure 6: Performance of continuous facial landmark tracking for each participant.
-FLAME 3D Head Model.
-IMUFace uses the FLAME [40] model to represent the face in 3D.
-FLAME is a statistical head model that learns from approximately 33,000 3D scans.
-It combines a PCA space for identity shape, simple rotational degrees of freedom for the neck, jaw, and eyeballs, pose-dependent corrective blend shapes, and global expression blend shapes to efficiently represent complex facial geometry and expressions.
-Mathematically, the model is defined as:
+4.4 3D Reconstruction
+To enhance system usability and simplify the modeling process, we aim to use a compact head model that balances reduced complexity with sufficient detail to produce expressive facial animations.
+The IMUFace system adopts the FLAME model (2020 version) for 3D facial representation.
+Specifically, the 3D reconstruction is primarily carried out using the FLAME PyTorch framework.
+FLAME is a statistical head model built from a dataset of approximately 33,000 3D scans.
+It efficiently captures complex facial geometries and expressions through a combination of identity PCA space, rotational degrees of freedom for the neck, jaw, and eyes, pose-dependent corrective blend shapes, and global expression blend shapes.
+The model is mathematically represented as (Eq.
+(8)):
 $FLAME(\vec{\beta},\vec{\theta},\vec{\psi}):\mathbb{R}^{|\vec{\beta}|\times|\vec{\theta}|\times|\vec{\psi}|}\to\mathbb{R}^{3N}$ (8)
-Optimization-based Shape Reconstruction.
-We employ a series of optimization steps to obtain a 3D face model that closely resembles the user’s real expressions.
-First, we perform camera calibration by optimizing the parameters for scale, rotation, and translation to minimize the L2 distance between the 2D projections of the 3D face model’s facial landmarks and the user’s facial landmarks.
-Second, we optimize the FLAME model parameters (shape, pose, and expression) to further reduce the L2 distance between the 2D projections of the facial landmarks from the 3D face model and the user’s facial landmarks.
-Ultimately, we can render the user’s 3D face model using the optimal FLAME model parameters.
-V Performance Evaluation
+A series of optimization steps is implemented to reconstruct a 3D face model that accurately reflects the user’s expressions.
+The process begins with camera calibration, where the parameters for scale, rotation, and translation are optimized to minimize the L2 distance between the 2D projections of the 3D model’s facial landmarks and the user’s facial landmarks.
+Following this, the FLAME model parameters—covering shape ($\vec{\beta}$), pose ($\vec{\theta}$), and expression ($\vec{\psi}$)—are fine-tuned to further reduce the L2 distance between the projected landmarks of the 3D model and the user’s facial landmarks.
+This L2-distance-based optimization method enables the model to fully adapt and reconstruct a unique shape that closely matches the target user’s facial features.
+Using the optimized FLAME parameters, the user’s 3D face model can then be rendered with high accuracy.
+While we customize the shape parameters for each user to capture their unique head geometry, we uniformly apply FLAME’s default texture parameters.
+This decision was primarily driven by our application’s focus on the dynamic representation of facial expressions and the accuracy of geometric shapes.
+The default texture provides a sufficient basic visual effect while avoiding the additional complexity and data acquisition costs associated with obtaining and fitting individual textures.
+Furthermore, a unified texture simplifies the rendering process, contributing to the system’s consistency and robustness.
+5 Performance Evaluation
 To validate the effectiveness of our system, we conducted extensive and detailed user studies to collect training data and evaluate the performance of our system in facial reconstruction.
@@
 These tests were designed to examine the system’s performance across various scenarios.
-V-A Experimental Methodology
-V-A1 Data Collection
+5.1 Experimental Methodology
+5.1.1 Data Collection
 In the study, we recruited 12 participants to evaluate IMUFace’s performance.
-The participants included nine males and three females from the local institution, aged from 18 to 28.
+The participants included nine males and three females from the local institution, aged 18 to 28.
 The experimental equipment comprises the IMUFace prototype, a Hikvision camera, and a desktop computer.
-Participants were asked to wear the IMUFace wireless earphones prototype, which was connected to the computer via Bluetooth.
+Participants were asked to wear the prototype of the IMUFace wireless earphones, which were connected to the computer via Bluetooth.
+This study was reviewed and approved by the Ethics Review Committee (ERC) at the South China University of Technology.
 To evaluate the performance of tracking 51 facial landmarks, we focused on seven universal facial expressions of emotion involving happiness, sadness, anger, surprise, fear, disgust, and contempt.
@@
 Each expression was made twice.
-Each data collection session lasted for 60 seconds, and the participants took adequate rest between sessions.
+Each data collection session lasted 60 seconds, and the participants took adequate rest between sessions.
 The earphone prototype was removed and remounted across sessions to validate the robustness to natural changes in sensor positions during daily usage.
-V-A2 Experimental Settings
+5.1.2 Experimental Settings
 The training was performed on a desktop with an Intel i9-12700K CPU, 32 GB of RAM, and an Nvidia GTX 4090 GPU.
@@
 To avoid overfitting issues that may happen in the training process, we added dropouts with a parameter of 0.1 following each ReLU activation.
-V-A3 Evaluation Metric
-The IMUFace system is evaluated primarily by two metrics [19]:
+5.1.3 Evaluation Metric
+The IMUFace system is evaluated primarily by two metrics [wu2021bioface]:
 1) Mean Absolute Error (MAE): the absolute difference between the predicted facial landmarks and the ground-truth facial landmarks.
@@
 $d_{norm}$} denotes the outer corner distance in the normalized coordinate frame.
-From Section IV-B3, we know that ${d_{norm}}=1$.
+From Section 4.2.3, we know that ${d_{norm}}=1$.
 This provided the MAE in units of millimeters, which is the primary metric used to evaluate IMUFace.
@@
 $d_{norm}$} denotes the outer corner distance in the normalized coordinate frame.
-From Section IV-B3, we know that ${d_{norm}}=1$.
-Figure 7: Domain adaptation.
-TABLE I: Comparison result of different Models.
-Models Average MAE Average NME
-Bi-LSTM [41] 2.57 mm 3.97%
-TCN [42] 2.39 mm 3.68%
-IF-ConvTransformer [43] 2.27 mm 3.50%
-Mamba2 [44] 2.37 mm 3.64%
-IMUTwinTrans 2.21 mm 3.40%
-V-B User-Dependent Evaluation
-For each user’s 12 sessions data, we randomly selected five sessions for training.
-The rest was used for testing in a randomized cross-validation fashion.
+From Section 4.2.3, we know that ${d_{norm}}=1$.
+Figure 6: Performance of continuous facial landmark tracking for each participant. (a) Per-participant Landmark Tracking Error. (b) MAE CDF for each participant. (c) Visualization of average MAE.
+5.2 User-Dependent Evaluation
+For each user’s 12 recorded sessions, we randomly selected five sessions for training, while the remaining sessions were used for testing in a randomized cross-validation manner.
 Fig.
-6(a) illustrates the average MAE & NME and corresponding standard deviations for all the 51 facial landmarks of each participant.
-Specifically, IMUFace obtains an average of 2.21 mm MAE and 3.40% NME with average standard deviations of 0.99 mm and 0.90%.
-Among all participants, U4 performed the best, with an MAE of only 1.46, while U8 had the highest MAE at 2.72.
+6(a) presents the average Mean Absolute Error (MAE) and Normalized Mean Error (NME), along with their corresponding standard deviations, for all 51 facial landmarks of each participant.
+Specifically, the blue bars represent the average MAE (in millimeters), which reflects the absolute spatial deviation between the predicted and ground truth landmarks.
+The red bars represent the average NME (in percentage), which measures the relative deviation normalized by the facial size, enabling cross-user performance comparison regardless of individual facial dimensions.
+These two complementary metrics provide a comprehensive evaluation of both absolute and relative reconstruction accuracy.
+IMUFace achieves an average MAE of 2.21 mm and an NME of 3.40%, with standard deviations of 0.99 mm and 0.90%, respectively.
+Among all participants, U4 achieves the best performance with an MAE of only 1.46 mm, while U8 exhibits the highest MAE at 2.72 mm.
 Fig.
-6(b) depicts the Cumulative Density Function (CDF) of the MAE errors for each individual participant as well as cross-participant cases.
-60.0% of the MAEs are less than 2.21, indicating a high level of prediction accuracy.
-Moreover, different landmarks may exhibit varying error levels due to their differing movement characteristics.
+6(b) illustrates the Cumulative Density Function (CDF) of MAE for each individual participant as well as across all participants.
+The CDF curve reveals that 60.0% of the samples have an MAE below 2.21 mm, indicating a high overall prediction accuracy.
+This distribution further confirms the robustness of our model in reconstructing facial landmarks with low error rates across different users.
+In addition, different landmarks tend to exhibit varying error levels due to their distinct movement patterns during facial expressions.
 Fig.
-6(c) visualizes the average MAE for all 51 major landmarks.
-We found that the reconstructed landmarks on the eyebrows and mouth have relatively more significant errors, while the facial landmarks around the eyes have more minor errors.
-This is mainly because the eyebrows and mouth change more during facial expression variations.
-V-C User-Adaptive Evaluation
+6(c) visualizes the average MAE for each of the 51 facial landmarks.
+We observe that landmarks located on the eyebrows and mouth tend to have higher reconstruction errors, while those around the eyes generally exhibit lower errors.
+This is mainly because the eyebrows and mouth undergo more dynamic deformations during expression changes, making their accurate reconstruction more challenging.
+5.3 User-Adaptive Evaluation
 Due to the need for improvement in cross-user testing results, we attempted to introduce domain adaptation techniques.
@@
 Through fine-tuning, the model can learn specific patterns from the target user’s data, thereby enhancing its prediction performance.
-V-D Micro-benchmark Tests
-V-D1 Model Selection
+Figure 7: Domain adaptation.
+Table 1: Comparison result of different Models.
+Models Average MAE Average NME
+Bi-LSTM [zhang2015bidirectional] 2.57 mm 3.97%
+TCN [verma2021expressear] 2.39 mm 3.68%
+IF-ConvTransformer [zhang2022if] 2.27 mm 3.50%
+Mamba2 [dao2024transformers] 2.37 mm 3.64%
+IMUTwinTrans 2.21 mm 3.40%
+Figure 8: Number of earbuds.
+5.4 Micro-benchmark Tests
+5.4.1 Model Selection
 To verify the effectiveness of our model, we compared it with several deep learning network methods commonly used for IMU data analysis to demonstrate the superiority of our model.
-We evaluated the performance of different networks under user-dependent experimental settings, and the results are presented in Table I.
-Figure 8: Number of earbuds.
+We evaluated the performance of different networks under user-dependent experimental settings, and the results are presented in Table 1.
 Figure 9: Training size vs error.
-V-D2 Number of Earphones
+5.4.2 Number of Earphones
 To validate the system’s efficacy under single-earphone conditions, we investigated the accuracy when using only a single earphone.
@@
 8, although using both earphones can provide the best accuracy by integrating sensor data from both ears, the accuracy with individual earphones (left or right) is also very close (2.40 mm, 2.36 mm).
-V-D3 Impact of Training Data Size
+5.4.3 Impact of Training Data Size
 In practical applications, the amount of data that needs to be collected from users for training can affect user experience.
@@
 It is shown that IMUFace can still achieve a reasonable level of accuracy even with a relatively short period of user data collection, which offers a high level of user-friendliness.
-V-D4 Latency and Power Consumption
+While the average MAE between two and five minutes of training data appears numerically close (2.27 mm vs.
+2.21 mm), a more detailed analysis reveals a noticeable improvement in the normalized mean error (NME), which decreases from 3.65% to 3.40%.
+Since NME accounts for differences in face scale and user variation, this reduction suggests that the model trained with five minutes of data has better generalization across diverse users and face shapes.
+This improvement in normalized prediction accuracy justifies the use of a longer training duration.
+Additionally, the extended data collection enhances the model’s robustness and stability in handling complex facial dynamics, making it more reliable for real-world applications.
+5.4.4 Comparison with Existing Solutions
+To compare the performance of IMUFace with existing vision-based and wearable methods, Table 2 presents the NME results across a range of representative systems.
+SDM [zhu2015face], CFSS [xiong2013supervised], and HRNet [wang2020deep] are classic vision-based approaches that rely on public image datasets such as 300-W to track 68 facial landmarks.
+In contrast, mm3DFace [xie2023mm3dface] and BioFace-3D [wu2021bioface] adopt alternative sensing modalities—millimeter-wave and EMG signals respectively—on self-collected datasets.
+EARFace [zhang2023earphone] is another non-visual method that uses in-ear sensing by emitting and receiving acoustic signals in the ear canal.
+Our method, IMUFace, achieves an NME of 3.40% on a self-collected dataset with 51 landmarks, which is very close to the performance of BioFace-3D (3.38%).
+However, unlike BioFace-3D, which requires attaching electrodes directly to the user’s face, IMUFace relies only on lightweight wearable IMU sensors, making it significantly more comfortable and user-friendly.
+Although IMUFace’s accuracy is slightly lower than EARFace (3.14%), it avoids several practical limitations of the EARFace system.
+Specifically, EARFace relies on wired earphones and uses active acoustic signal emission into the ear canal, which may pose risks to user hearing health, involves higher power consumption, and compromises convenience due to its tethered design.
+In contrast, IMUFace offers comparable performance while being low-power, wireless, and non-invasive.
+These advantages highlight its strong potential for practical deployment, particularly in scenarios where visual input is unavailable or privacy-sensitive environments are involved.
+Table 2: Comparison with existing solutions.
+Methods Dataset Landmarks NME (%)
+SDM [zhu2015face] 300-W 68 7.52
+CFSS [xiong2013supervised] 300-W 68 5.76
+HRNet [wang2020deep] 300-W 68 2.87
+mm3DFace [xie2023mm3dface] Self-collected 68 3.94
+BioFace-3D [wu2021bioface] Self-collected 53 3.38
+EARFace [zhang2023earphone] Self-collected 51 3.14
+IMUFace (Ours) Self-collected 51 3.40
+5.4.5 Latency and Power Consumption
 The inference time for the 51 facial landmarks was measured on a single NVIDIA GTX 4090 GPU.
@@
 This indicates that with a 100 mAh lithium polymer battery, IMUFace can continuously record data for up to 6.5 hours, meeting the needs of most applications.
-We examined the power consumption of recent studies on 3D facial reconstruction and summarized the overall power consumption of the system, including the use of sensing and communication power, in Table II.
-Our analysis shows that IMUFace consumes the least power among all sensing technologies evaluated, with energy consumption 58.0% lower than that of the next most efficient technology, Bioface-3D [19].
-TABLE II: Comparison of System Power Consumption among Related Existing Studies
+We examined the power consumption of recent studies on 3D facial reconstruction and summarized the overall power consumption of the system, including the use of sensing and communication power, in Table 3.
+Our analysis shows that IMUFace consumes the least power among all sensing technologies evaluated, with energy consumption 58.0% lower than that of the next most efficient technology, Bioface-3D [wu2021bioface].
+Table 3: Comparison of System Power Consumption among Related Existing Studies
 Studies Sensors Power Consumption
-NeckFace [12] Cameras 4 W
-C-Face [22] Cameras >4 W
-BioFace-3D [19] Biosensors 138 mW
-EarIO [20] Acoustics 153.7 mW
-EyeEcho [15] Acoustics 167 mW
+NeckFace [chen2021neckface] Cameras 4 W
+C-Face [chen2020c] Cameras >4 W
+BioFace-3D [wu2021bioface] Biosensors 138 mW
+EarIO [li2022eario] Acoustics 153.7 mW
+EyeEcho [li2024eyeecho] Acoustics 167 mW
 IMUFace IMUs 58 mW
-VI Discussion
-Our study aims to use conventional motion sensors, such as accelerometers and gyroscopes, to collect data.
-These sensors are practical and portable for use in any environment.
-However, the facial expression movement characteristics of users differ significantly, leading to noticeable variations in model performance across different users.
-Therefore, the system requires a small amount of data from the target user.
-VI-A User-independent Model.
-Through the experimental results, we found that the performance of the model after fine-tuning with user data is significantly better than that of the user-independent model.
-This is because the intensity, response, and sensitivity of facial movement data can vary significantly from person to person, leading to substantial differences between data domains.
-However, this may reduce usability, as new users must go through a registration phase before using the system.
-Our future research will explore generating large amounts of virtual IMU data using data from other modalities to train a more generalized user-independent model.
-VI-B Effects of Body Movements.
-In the system design, we considered the impact of initial head pose changes on the IMU signals.
-However, repetitive head bounces during activities such as walking can produce significant inertial signal noise [45], which can interfere with the normal operation of the system.
-Future research can explore the use of targeted filtering methods to minimize the noise impact introduced by such activities.
-VII Conclusion
+6 Discussion
+This study explores the use of conventional motion sensors, such as accelerometers and gyroscopes, for capturing facial expression data, emphasizing their practicality and portability for a wide range of applications.
+A key challenge identified is the variability in facial movements across different users, which impacts the model’s performance.
+To address this, incorporating a small amount of user-specific data has been shown to improve accuracy.
+This approach offers promising applications in fields like healthcare, gaming, and human-computer interaction by advancing real-time facial expression tracking and enabling innovative functionalities.
+6.1 User-independent Model
+Experimental results reveal that fine-tuning the model with user-specific data significantly enhances performance compared to a purely user-independent model.
+This improvement arises from the inherent variability in facial movement characteristics, such as intensity, responsiveness, and sensitivity, among individuals.
+These differences create unique data domains, posing challenges for generalized models.
+Although fine-tuning improves accuracy, it introduces a trade-off in usability, as new users must register and provide data for personalization.
+To address this limitation, future work will focus on generating extensive virtual IMU datasets using multimodal inputs, such as video or depth sensors.
+By leveraging synthetic data, we aim to train a robust and generalized user-independent model capable of adapting to a broader user base with minimal personalization.
+Additionally, investigating techniques like domain adaptation and transfer learning could help bridge domain gaps, enhancing the model’s generalization capabilities.
+These strategies aim to maintain high performance while improving the system’s usability for diverse users.
+6.2 Effects of Body Movements
+While the system accounts for variations in initial head pose, dynamic body movements—such as repetitive head bounces during activities like walking or running—introduce significant noise into IMU signals [prakash2019stear].
+This noise complicates the accurate detection of facial expressions, reducing the system’s reliability in real-world scenarios.
+To mitigate this issue, future research will explore advanced filtering techniques designed to isolate and remove noise caused by body movements while preserving subtle facial expression signals.
+Methods such as adaptive filtering, wavelet transforms, or machine-learning-based noise suppression may improve signal quality.
+Furthermore, integrating complementary sensors, such as magnetometers or pressure sensors, could provide additional context to distinguish between facial and body-induced movements.
+These enhancements aim to ensure reliable system performance in dynamic and mobile environments.
+6.3 Practical Applications and Industry Implications
+The proposed methodology demonstrates significant potential across various sectors, highlighting its practical value.
+In healthcare, it enables remote monitoring of patients with neurological disorders, facilitating the assessment of facial muscle activity and therapy progress.
+In gaming and virtual reality, it offers camera-free, real-time facial expression tracking, enhancing user privacy and reducing hardware complexity.
+For human-computer interaction, it supports more intuitive communication between users and devices.
+In the automotive industry, the system can monitor driver fatigue or emotional states, improving safety and user experience.
+These diverse applications underline the method’s versatility and its potential to address practical challenges across multiple disciplines.
+7 Conclusion
 This article proposes and designs IMUFace, a system that reconstructs the 3D facial expressions of users using IMUs that are attached to (and potentially built into) wireless earphones.